Question 3 0 out of 10 points A closed rigid still vessel contains a gas. The initial minus final change in internal energy of the system equals 90J.84J of heat are lost to the surroundings. Calculate the work exchanged during the process, Give the answer with zero decimal places. Selected Answer:6 Correct Answer: 6+ Response Feedback: Please, try again
Answer #1

Given data,

Initial minus final change in internal energy = U1 - U2 = 90J

Heat lost to the surrounding = Q = -84J (Heat rejected from system)

Assumptions

accrding to question, it is rigid still vessel

therefore, change in kinetic energy = KE = 0.

Therefore, Change in potential energy = PE = 0.

Considering gas in the tank as a system,

tank is rigid theterfeo, no boundary work.

Applyinng energy balance on the system.

Net energy transfer by heat and work = chnage in internal energy.

Therefore, work exhange during process = W = -6J (Answer ).

Give data ange in internal energy Heat lost to the scr roundingニG- Ass CK ange ir> kioenc energy = AKE-O . change in potential energy-APE-o Sao Systern sideing ao in -the tank as a Syefem Appltig energy balarce on he Syesterm Net energy tansfer by heatand wosk- charge in indtesna) energ -3 hơrk exchange duoi ng process = W--6J (Ang)